Calculate the electric field strength from a 1.25×10-8 C charge if the test charge is placed 2.50×10-5m.

1
Expert's answer
2022-05-18T08:31:12-0400
"E=k\frac{q}{r^2}"

"E=9*10^9*\frac{1.25*10^{-8}}{(2.5*10^{-5})^2}=1.8*10^{11}\:\rm N/C"